Sipur and Mispar in Ibn Ezra’s Sefer HaMispar Doron ZEILBERGER^{1}

Rabbi Abraham Ben Meir Ibn Ezra (1092-1167) was one of the all-time greatest in the following categories:

• Poetry

• Biblical Exegesis

• Grammar

• Philosophy

• Mathematics But he was most proud of his contributions to

• Astrology He had a very hard life, as he testifies himself in one of his many poems (translated by Leon J. Weinberger[W], p. 65) Early I set out for the patron’s home; They say: He is off riding; I return toward evening, They say: He is already sleeping; He either mounts a horse or climbs into bed; Woe to the wretched man who is unlucky! But Ibn Ezra took it well, since he knew that he was ‘born without a star’, and that was his destiny. Unlike other Jewish notables in Spain of the time, who were well-to-do, many of whom making their living as doctors, our hero was usually penny-less. There were two particularly tragic event in his life. The first one being exiled by a new regime of ‘fundamentalist’ Moslems, who displaced the former enlightened rule of moderate Islam, that gave the Jews the alternative to convert or leave. That’s how he winded up in Italy, France, and England, and perhaps it was just as well, because all of his important books and biblical commentary were written then, in part to support himself. A bigger tragedy was that his son Isaac, a notable poet on his own right, converted to Islam. This inspired a beautiful eulogy ([W], p. 92):

1 Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen

Rd., Piscataway, NJ 08854-8019, USA. zeilberg at math dot rutgers dot edu , http://www.math.rutgers.edu/~zeilberg . First version: Oct. 14, 2005. Transcript of a talk delivered (July 12, 2005, 10:30-11:00am local time) at the “Math and Narrative” meeting organized by Thales & Friends under the leadership of Apostolos Doxiadis, Mykonos, July 12-15, 2005. Supported in part by the NSF.

Father to child, approach and mourn

For God has removed from you

Your son, your only son

Whom you loved, even Issac

Ibn Ezra was way ahead of his time in many ways. For example, his abstract, pantheist notion of God, that deeply influenced Baruch Spinoza five hundred years later. He was also a pioneer in his attitude to sex. In his commentary to Leviticus he says:

It is known that there are three kinds of sexual intercourse, the first for procreation without desire, the second for easing the weariness of the body and the third for satisfying beastly desire

So he extended the traditional dichomoty of good vs. bad sex to a trichotomy, recoginzing that sometimes we need to have sex for mental and physical health reasons.

Ibn Ezra was a also a great puzzlist. The following version of the extended Josephus problem is attributed to him.

One day the rabbi was in a ship with fifteen of his disciples and fifteen crooks, altogether thirty people. And it came to pass that God sent a big storm, and the sea was so high that the ship was about passages of his famous biblical commentary. But he only wrote one mathematical text, “Sefer HaMispar” which means “book of number”. I wish that all textbooks would be written like this one. First, it is a pedagogical masterpiece, since